OpenStudy (august899):

Determine which expression is equivalent to cos3 θ. (Multiple Choice)

OpenStudy (august899):

A. 3sin^3 θ-4cos θ B.3cos^4 θ-4cos θ C.4cos^3 θ-3cos θ D.4sin^3 θ-3cos θ

OpenStudy (mww):

There's two ways you can do this one. 1) by the expansion cos (x+y) = cosx cosy - sinx siny 2) by equating De Moivre's Theorem with binomial theorem terms The way most people do is method one: \[\cos(3 \theta) = \cos(2\theta + \theta) = \cos(2 \theta)\cos(\theta) - \sin(2\theta) \sin(\theta)\] Then substitute the cos(2 theta) and sin (2theta) double angle identities in and see what you get

OpenStudy (august899):

So is it like this? cos(2θ)=cos(1θ+1θ)=cos(1θ)cos(1θ)-sin(1θ)sin(1θ)

OpenStudy (mww):

keep going. Simplify further

OpenStudy (august899):

I don't know what to do after that.

OpenStudy (mww):

you should get \[\cos(2 \theta) = \cos^2(\theta) - \sin^2(\theta) = 2\cos^2 (\theta) - 1 = 1 - 2 \sin^2(\theta)\] (the three are interchangeable) And \[\sin(2 \theta) = 2\sin(\theta)\cos(\theta)\]

OpenStudy (august899):

Ok. I think it's C.

OpenStudy (mww):

then sub these two back into the orginal \[cos(3 theta) = cos(2 theta + theta) = cos(2 theta) cos(theta) - sin(2theta) sin(theta)\] \[= (2\cos^2\theta -1)\cos \theta - 2\sin \theta \cos \theta(\sin \theta) = 2\cos^3(\theta) - \cos \theta -2\sin^2(\theta)\cos(\theta)\] When you get sin^2(theta), replace it with 1 - cos^2(theta) (Pythagorean identity) \[2\cos^3\theta - \cos \theta - 2(1-\cos^2(\theta))\cos(\theta) = 2\cos^3 \theta - \cos \theta -2\cos \theta + 2\cos^3 \theta\] \[= 4\cos^3 \theta - 3 \cos \theta\]

OpenStudy (killuaxzoldyck):

https://akk.li/pics/anne/jpg